A volume of 400 cc of air is measured at a pressure of 740 mm Hg abs and a temperature of 18°C. what will be the volume at 760 mm Hg abs and 0°C?  


1
Expert's answer
2021-10-18T11:01:54-0400

"V_1 = 400 \;cc \\

p_1 = 740 \;mmHg \\

T_1 = 18 +273 = 291 \;K \\

p_2 = 760 \;mmHg \\

T_2 = 0 +273 = 273 \;K \\

\frac{p_1V_1}{T_1} = \frac{p_2V_2}{T_2} \\

V_2 = \frac{p_1V_1T_2}{T_1p_2} \\

V_2 = \frac{740 \times 400 \times 273}{291 \times 760} = 365.4 \;cc"

Answer: 365.4 cc